OpenStudy (anonymous):

Find the points of inflection for the curves; y = 5 - 2x - 3x² - x³

OpenStudy (anonymous):

what's inflection?? :P

OpenStudy (anonymous):

It's like the changes in curve and the sense of concavity.. :)

OpenStudy (anonymous):

ohhhhhhh. so you mean the points where the direction changes. i feel stupid now that i don't know what this is :P

OpenStudy (anonymous):

(-1,5)

myininaya (myininaya):

y'=0-2-6x-3x^2 y''=0-0-6-6x=-6-6x set y''=0 and solve for x -6-6x=0 x=-1 if x=-1 then y=5-2(-1)-3(-1)^2-(-1)^3=5+2-3+1=5 does the inflection change at (-1,5) well lets test the concavity before and after x=-1 choose a number before -1 such as 0; f''(0)=-6 choose a number after -1 such as 2; f''(2)=-6-6(2)=-6-12=-18 both are negative that means the concavity did not change that means there are no inflection points

OpenStudy (amistre64):

f' gives slope f'' gives concavity f' = 0 gives min max f'' = 0 gives inflection

myininaya (myininaya):

i mean 0 isn't before -1

myininaya (myininaya):

lol

OpenStudy (amistre64):

which is it? ahhhhh!!!!!

myininaya (myininaya):

choose a number before -1 such as -2; f''(-2)=-6-6(-2)=6 concavity changes=> inflection point at (-1,5)

myininaya (myininaya):

i'm gonna go write my 1,2,3,'s now later

OpenStudy (anonymous):

I get it.. (±2, 5) thanks!

OpenStudy (amistre64):

..... no

OpenStudy (amistre64):

a cubic only has one inflection

myininaya (myininaya):

y'=0-2-6x-3x^2 y''=0-0-6-6x=-6-6x

OpenStudy (anonymous):

Oh alright..

OpenStudy (amistre64):



OpenStudy (anonymous):

art is really improving. like in the renaissance

myininaya (myininaya):

lol

OpenStudy (anonymous):

amsitre got it

OpenStudy (anonymous):

now i am out before i get stuck